An electric kettle is connected to a 120-V

outlet and has 11.5 A of current flowing

through it. What is the resistance in the

kettle?


1
Expert's answer
2022-06-08T08:07:03-0400

Given:

"U=120V;"

"I=11.5A;"

Find:

"R" - resistance of the kettle.

Solution:

Let's use Ohm's law fro part of circuit:

"I=\frac{U}{R}" , "R=\frac{U}{I}=\frac{120V}{11.5A}=10.435\Omega."

Answer:

"R=10.435\Omega."